Assumes Qd = 1000 – 10P Qs = 10P

If a 20$ subsidy is placed on the Good What is new Pc, Pp, and Q.

What is Consumer Surplus?

What Is Producer Surplus?

What is the cost to the government?

What is Total Surplus?

What is dead weight loss?

What is Total Surplus?

What is dead weight loss?
Answer #1

Subsidy= $20

New Supply Equation

Qs'= 10(P+20)=10P+200

Equilibrium

10P+200=1000-10P

20P= 800

P*= $40

Q*= 10*40+200=600

Price paid by consumers=Pc= $40

Price received by producer= Pp=$40+$20=$60

Equilibrium quantity=Q= 600

Consumer surplus= 0.5(100-40)(600)= $18000

Producer surplus=0.5(60)(600)=$18000

Cost to the government= Subsidy *Quantity=$20*600=$12000

Deadweight loss= 0.5(20)(600-500)= $1000

Total surplus=$18000+$18000-$12000= $24000